OpenStudy (anonymous):

find all numbers for which the rational expression is undefined (t^3-2t)/(t^2-49)

jhonyy9 (jhonyy9):

t(t*2 -2)=0 --- t_1=0 --- t*2 = 2 --- t_2,_3 =+/- radical 2 t*2 -49 =0 t*2 =49 t_4,_5 = +/- 7 - so for those solutions will be undefined this expression .